# Hume: Introducing System-2 Thinking in Visual-Language-Action Model

### v1

@@ 5 Conclusion and Limitations
[…]
Supplementary Material
6 Hume Visualization Analysis
In this section, we first demonstrate the detailed inference process of Hume through the dexterous Push-T task in Sec. 6.1, then visualize two key designs of Hume: value-guided System-2 thinking and cascaded action denoising in Sec. 6.2 and Sec. 6.3 to provide a comprehensive understanding of Hume.
6.1 Hume Workflow Visualization
@@
6.2 Value-Guided Thinking Visualization
Fig. 7 visualizes estimated state-action values the candidate actions at different time steps in the LIBERO-GOAL setting. Since the action space in LIBERO is 7-dimensional, we used Principal Component Analysis (PCA) to project the high dimensional actions onto a two-dimensional plane. For each projected candidate action, we use different colors to represent their corresponding state-action values estimated by the value-query head. These projected points together with their corresponding state-action values generate the value map shown in the figure. In the value map, yellow color represent actions with high state-action values, while purple color represent actions with low state-action values. Additionally, we also show the ground truth actions from collected demonstrations in the value map for comparison.
Figure 7: Value Map of Candidate Actions. The candidate actions $\mathbf{A}_{t}^{\tau_{i}}$ sampled from System 2 and ground-truth actions $\mathbf{A}_{t}^{\text{GT}}$ are projected into the same two-dimensional space through Principal Component Analysis (PCA). The intensity of colors indicates the magnitude of state-action values $Q(\mathbf{q_{t}},\mathbf{A}_{t}^{\tau_{i}})$ of candidate actions.
@@ 6.2 Value-Guided Thinking Visualization
6.3 Cascaded Action Denoising Visualization
@@
7 Video Demonstration and Anonymous Link
We provide a video of Hume and an anonymous link (Please refer to: https://hume-vla.github.io for more details) to demonstrate the deployment on real-world robot platforms.
8 Implementation Details
@@
8.1 Value Objective Functions
@@
8.2 Training and Inference Hyperparameters
@@
9 Experiment Details
In this section, we will provide experiment details including evaluation setup and additional test results. In Sec. 9.1, we provide detailed descriptions of test setup for standard simulation benchmark and the implementation details of all comparison methods along with the detailed test results. In Sec. 9.2, we introduce the detailed setup of test tasks on real robot platforms and testing standards, and provide the detailed test results.
9.1 Simulation Benchmark Details
@@
9.2 Real-World Evaluation Details

### v2

@@ 5 Conclusion and Limitations
Appendix A Hume Visualization Analysis
In this section, we first demonstrate the detailed inference process of Hume through the dexterous Push-T task in Sec. A.1, then visualize two key designs of Hume: value-guided System-2 thinking and cascaded action denoising in Sec. A.2 and Sec. A.3 to provide a comprehensive understanding of Hume.
A.1 Hume Workflow Visualization
@@
A.2 Value-Guided Thinking Visualization
Fig. 2 visualizes estimated state-action values of the candidate actions at different time steps in the LIBERO-GOAL benchmark. Since the action space in LIBERO is 7-dimensional, we used Principal Component Analysis (PCA) to project the high dimensional actions onto a two-dimensional plane. For each projected candidate action, we use different colors to represent their corresponding state-action values estimated by the value-query head. These projected points together with their corresponding state-action values generate the value map shown in the figure. In the value map, yellow color represent actions with high state-action values, while purple color represent actions with low state-action values. Additionally, we also show the ground truth actions from collected demonstrations in the value map for comparison.
@@ A.2 Value-Guided Thinking Visualization
A.3 Cascaded Action Denoising Visualization
@@
Appendix B Implementation Details
@@
B.1 Value Objective Functions
@@
B.2 Training and Inference Hyperparameters
@@
Appendix C Experiment Details
In this section, we will provide experiment details including evaluation setup and additional test results. In Sec. C.1, we provide detailed descriptions of test setup for standard simulation benchmark and the implementation details of all comparison methods along with the detailed test results. In Sec. C.2, we introduce the detailed setup of test tasks on real robot platforms and testing standards, and provide the detailed test results.
C.1 Simulation Benchmark Details
@@
C.2 Real-World Evaluation Details